# Error Propagation Mechanisms and Compensation Strategies for Quantized Diffusion Models

### v2

@@ 4.2 Main Results
Image Generation Evaluation. As shown in Table 1, we conduct extensive experiments at two quantization precisions: W8A8 and W4A4. We observe that, across all precision levels, TCEC achieves better image fidelity and diversity, and it even matches the 16-bit results under W8A8 quantization. For UNet-based models, on SDXL, our W4A4 model substantially outperforms SVDQuant W4A4, the current SOTA 4-bit approach, achieving an sDCI PSNR of 21.9. This even surpasses TensorRT’s W8A8 result of 21.7, demonstrating robust performance under lower-bit quantization. On SDXL-Turbo, MixDQ W4A4 exhibits abnormal FID and IR metrics, indicating quantization failure. This highlights the greater difficulty of quantizing models with a small number of inference steps. Our W4A4 model surpasses SVDQuant by 0.017 and 0.02 in MJHQ IR and sDCI IR metrics, respectively, suggesting a stronger alignment with human preferences. The larger performance gains of TCEC on SDXL than SDXL-Turbo highlight the critical role of cumulative error correction in longer inference sequences. For DiT-based model, on PixArt-$\sum$, our W4A4 model significantly surpasses SVDQuant’s W4A4 results across all metrics. Leveraging the DPM++ solver, PixArt-$\sum$ demonstrates TCEC’s robustness across different solver configurations. As shown in Figure 2, when compared with ViDiT-Q W4A4 and SVDQuant W4A4, the TECE W4A4 method demonstrates less quality degradation and smaller changes in image content. More visual results can be found in Appendix H.
Video Generation Evaluation. As shown in Table 2, our evaluation compares SOTA PTQ methods under both W8A8 and W4A8 configurations. In Imaging Quality, TCEC achieves 65.56 (W8A8) and 64.97 (W4A8), surpassing ViDiT-Q by +2.08 and +3.90 absolute points respectively. It demonstrates significant advantages in frame-wise quality metrics that evaluate static visual fidelity independent of temporal factors. TCEC achieves 96.27 motion smoothness at 8-bit and 96.01 at 4-bit quantization, outperforming ViDiT-Q by 0.13 and 0.32 respectively. This validates the effectiveness of our temporal-channel decoupled compensation strategy in handling error accumulation across denoising steps. The scene consistency metric reaches 39.58 (8-bit) and 37.20 (4-bit), establishing 1.36 and 1.01 improvements over baselines, which confirms stable long-sequence generation through temporal error propagation modeling.
Hardware Resource Savings. TCEC constructs its quantization strategy based on SVDQuant by integrating cumulative error correction mechanisms. As TCEC’s research focus is not on engineering performance, we fully reuse its inference engine, Nunchaku (Li et al., 2024b). As shown in Table 3, experimental tests indicate that the resulting end-to-end (E2E) latency degradation is less than $0.5\%$. Further analysis can be found in the Appendix G. Therefore, on one 12B model, TCEC-W4A4 reduces the memory footprint by $3.5\times$ compared to the FP16 model and accelerates inference by $3\times$ versus NF4 weight-only quantization on the laptop-level RTX 4090, with engineering performance comparable to SVDQuant and significantly outperforming other PTQ algorithms.
@@
Table 2: Performance of PTQ Algorithms for OpenSora on Vbench eval Benchmark.In the W4A8 configuration, TCEC outperforms other methods.
Method Bit-width W/A Imaging Quality Aesthetic Quality Motion Smooth. Dynamic Degree BG. Consist. Subject Consist. Scene Consist. Overall Consist. - 16/16 63.68 57.12 96.28 56.94 96.13 90.28 39.61 26.21 Q-Diffusion 8/8 60.38 55.15 94.44 68.05 94.17 87.74 36.62 25.66 Q-DiT 8/8 60.35 55.80 93.64 68.05 94.70 86.94 32.34 26.09 PTQ4DiT 8/8 56.88 55.53 95.89 63.88 96.02 91.26 34.52 25.32 SmoothQuant 8/8 62.22 55.90 95.96 68.05 94.17 87.71 36.66 25.66 Quarot 8/8 60.14 53.21 94.98 66.21 95.03 85.35 35.65 25.43 ViDiT-Q 8/8 63.48 56.95 96.14 61.11 95.84 90.24 38.22 26.06 ViDiT-Q + TCEC 8/8 65.56 57.12 96.27 61.09 96.23 91.34 39.58 26.20 Q-DiT 4/8 23.30 29.61 97.89 4.166 97.02 91.51 0.00 4.985 PTQ4DiT 4/8 37.97 31.15 92.56 9.722 98.18 93.59 3.561 11.46 SmoothQuant 4/8 46.98 44.38 94.59 21.67 94.36 82.79 26.41 18.25 Quarot 4/8 44.25 43.78 92.57 66.21 94.25 84.55 28.43 18.43 ViDiT-Q 4/8 61.07 55.37 95.69 58.33 95.23 88.72 36.19 25.94 ViDiT-Q + TCEC 4/8 64.97 56.90 96.01 59.42 97.01 90.05 37.20 26.20
@@
Table 3: A comparison of inference overhead between SVDQuant and SVDQuant with TCEC.
Model FP16 SVDQuant W4A4 SVDQuant+TCEC W4A4
Flux-dev 1.0 (T = 30) 26.14 s 9.947 s 9.996 s
@@ 5 Conclusion
Appendix
Appendix A LLM Usage
Large Language Models were used solely to refine the manuscript’s language, including sentence rephrasing, grammar checking, and improving readability. The LLM was not involved in ideation, methodology, experimental design, or data analysis. All scientific content, concepts, and analyses were developed by the authors, who take full responsibility for the manuscript. The LLM’s role was limited to linguistic polishing, with strict adherence to ethical guidelines and avoidance of plagiarism or scientific misconduct.
@@ Appendix C Empirical Study of Error Accumulation Steps
In this section, we extend TCEC to other solvers to demonstrate its generality, such as the most commonly used high-order solver DPM++ (Lu et al., 2022). The iterative update for DPM-Solver++ (2nd-order variant) is given by:
@@ Appendix F Ablation experiment of regularization term $\lambda_{1}$
Appendix G Inference Overhead of TCEC
@@
TCEC improves performance by performing error correction on the output of the quantized model, and it is completely orthogonal to the backbone quantization algorithm. As shown in Eq. 14, the computation of single-step quantization error involves no complex operations and is accomplished solely through a non-linear mapping: $\varepsilon_{t}=\mathbf{K}_{t}\odot\tilde{\mu}_{\theta}(\tilde{x}_{t},t)$, where $\mathbf{K}_{t}$ denotes a timestep-conditioned channel-wise scaling matrix, and $\odot$ represents element-wise multiplication. Consequently, the additional theoretical computational complexity introduced at each step is $NC^{2}HW$, which is negligible compared to that of the DIT/Unet-Backbone. Practical test data on the Nvidia A800-40G platform show in Table 3 and Table 6 that the extra end-to-end latency incurred is less than 0.5%, with specific test data provided in the table below.
Table 6: A comparison of inference overhead between ViDiT-Q and ViDiT-Q with TCEC.
Model FP16 ViDiT-Q W8A8 ViDiT-Q+TCEC W8A8
Opensora1.2 (51 frames, 480P) 44.56 s 26.211s 26.316s
CogVideoX (48 frames, 480P) 78.48 49.67s 49.894s
Wan2.1-1.3B (81 frames, 480P) 199 s 118.45s 119.029s
Appendix H Visual Quality Results.

### v3

@@ 4.2 Main Results
Image Generation Evaluation. As shown in Table 1, we conduct extensive experiments at two quantization precisions: W8A8 and W4A4. We observe that, across all precision levels, TCEC achieves better image fidelity and diversity, and it even matches the 16-bit results under W8A8 quantization. For UNet-based models, on SDXL, our W4A4 model substantially outperforms SVDQuant W4A4, the current SOTA 4-bit approach, achieving an sDCI PSNR of 21.9. This even surpasses TensorRT’s W8A8 result of 21.7, demonstrating robust performance under lower-bit quantization. On SDXL-Turbo, MixDQ W4A4 exhibits abnormal FID and IR metrics, indicating quantization failure. This highlights the greater difficulty of quantizing models with a small number of inference steps. Our W4A4 model surpasses SVDQuant by 0.017 and 0.02 in MJHQ IR and sDCI IR metrics, respectively, suggesting a stronger alignment with human preferences. The larger performance gains of TCEC on SDXL compared to SDXL-Turbo highlight the critical role of cumulative error correction in longer inference sequences, and we provide a complete exploration of the minimal inference steps required for high-speed inference in Appendix G. For DiT-based model, on PixArt-$\sum$, our W4A4 model significantly surpasses SVDQuant’s W4A4 results across all metrics. Leveraging the DPM++ solver, PixArt-$\sum$ demonstrates TCEC’s robustness across different solver configurations. As shown in Figure 2, when compared with ViDiT-Q W4A4 and SVDQuant W4A4, the TECE W4A4 method demonstrates less quality degradation and smaller changes in image content. More visual results can be found in Appendix K.
@@
Table 2: Performance of PTQ Algorithms for OpenSora on Vbench eval Benchmark in the W4A8 configuration.
[…]
@@
Table 3: Inference overhead comparison with TCEC. Quantization: video models use W8A8, Flux-dev uses W4A4.
Model Type FP16 (s) Quantized Inference (s) w/o TCEC w/ TCEC Opensora1.2 (51 frames, 480P) Video (W8A8) 44.56 26.211 26.316 CogVideoX (48 frames, 480P) Video (W8A8) 78.48 49.67 49.894 Wan2.1-1.3B (81 frames, 480P) Video (W8A8) 199 118.45 119.029 Flux-dev 1.0 (T = 30) Img2Vid (W4A4) 26.14 9.947 9.996
Video Generation Evaluation. As shown in Table 2, our evaluation compares SOTA PTQ methods under both W8A8 and W4A8 configurations. In Imaging Quality, TCEC achieves 65.56 (W8A8) and 64.97 (W4A8), surpassing ViDiT-Q by +2.08 and +3.90 absolute points respectively. It demonstrates significant advantages in frame-wise quality metrics that evaluate static visual fidelity independent of temporal factors. TCEC achieves 96.27 motion smoothness at 8-bit and 96.01 at 4-bit quantization, outperforming ViDiT-Q by 0.13 and 0.32 respectively. This validates the effectiveness of our temporal-channel decoupled compensation strategy in handling error accumulation across denoising steps. The scene consistency metric reaches 39.58 (8-bit) and 37.20 (4-bit), establishing 1.36 and 1.01 improvements over baselines, which confirms stable long-sequence generation through temporal error propagation modeling. In addition, Table 9 in the Appendix reports a comparison between TCEC and QNCD, demonstrating the effectiveness of TCEC in cumulative error correction.
Hardware Resource Savings. TCEC builds its quantization strategy upon ViDiT-Q and SVDQuant by incorporating cumulative error correction mechanisms. Since TCEC does not aim to optimize engineering-level performance, we fully reuse the original ViDiT-Q and SVDQuant inference engine, to ensure a fair comparison. As shown in Table 3, rows 1–3 adopt ViDiT-Q quantization with the TCEC overlay, whereas row 4 applies SVDQuant quantization with the same TCEC overlay. Experimental results indicate that the additional end-to-end (E2E) latency overhead introduced by TCEC is less than $0.5\%$. More detailed analyses are provided in Appendix H. Consequently, on a 12B model, TCEC-W4A4 achieves a $3.5\times$ reduction in memory footprint compared to the FP16 baseline and delivers a $3\times$ inference speedup over NF4 weight-only quantization on a laptop-class RTX 4090, while maintaining engineering efficiency comparable to SVDQuant and significantly outperforming other PTQ methods.
@@ 5 Conclusion
Impact Statement
This paper aims to advance research on model compression in the field of generative models. Our work does not involve any personally identifiable information nor does it conduct experiments that may raise privacy or security concerns.
Appendix A Explanation of Approximation 1
In Approximation 1, we approximate $J_{x_{t}}\approx 0$. However, we clarify that this approximation does not ‘deny the existence of the Jacobian’. Rather, based on magnitude analysis of error propagation and the inherent properties of diffusion models, we argue that its contribution to cumulative error is negligible compared to the dominant terms. This makes the approximation a theoretically grounded and empirically supported simplification. The details are as follows: 1.Core rationale of the approximation: a magnitude-first simplification. In the error propagation equation (Eq. 7), the propagation matrix $A_{t}=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}I+B_{t}\cdot\mathbf{J}_{x_{t}}$ is composed of two components:
  • Dominant term $\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}I$ : This term arises from the diffusion noise-scheduling mechanism (where $\alpha_{t}$ is tied to the noise variance $\beta_{t}$). Its magnitude consistently falls in the 0.9–1.0 range, making it the primary driver of error propagation, typically contributing over 95% of the total effect.
  • Secondary term $B_{t}\cdot\mathbf{J}_{x_{t}}$: Here, $B_{t}$ is an $\alpha_{t}$-dependent coefficient with magnitude $\leq$ 0.05. Meanwhile, for a well-trained diffusion model optimized via “denoising score matching,” the noise estimator $\mu_{\theta}$ satisfies a Lipschitz continuity condition with constant $L<0.3$, implying $\|\mathbf{J}_{x_{t}}\|\leq L$. Consequently, the magnitude of $B_{t}\cdot\mathbf{J}_{x_{t}}$ is at most 1/100–1/10 of the dominant term. Its influence on the distribution of accumulated error is therefore minimal.
2. Support from both theoretical insights and empirical evidence. From a theoretical standpoint, the intermediate states $x_{t}$ in diffusion models are well known to approximate a standard Gaussian distribution at most timesteps due to strong noise injection. This heavy smoothing of the input space forces the model to focus primarily on global noise evolution rather than local perturbations, which naturally suppresses the influence of the Jacobian term $\mathbf{J}_{x_{t}}$. From an empirical standpoint, as show in Table 4 we measure the ratio between the secondary term and the dominant term across different timesteps using SDXL under W4A4 quantization. Even in low-noise regimes (e.g., timestep $t=40$), the secondary term contributes only 0.45% relative to the dominant term—well within the threshold of being safely negligible in magnitude.
Table 4: Statistics of dominant and secondary terms at different SDXL timesteps.
time step (SDXL, T=50) Dominant term $\mathbf{J}_{x_{t}}$ $B_{t}\cdot\mathbf{J}_{x_{t}}$ Secondary / Dominant
10 (High-noise stage) 0.982 0.021 0.00105 0.107%
25 (Intermediate stage) 0.951 0.037 0.00185 0.194%
40 (Low-noise stage) 0.915 0.083 0.00415 0.454%
3. Necessity of the simplification. If the Jacobian term were retained, computing the full error propagation would require matrix inversion and higher-order derivatives, causing the computational complexity to surge from $\mathcal{O}(T)$ to $\mathcal{O}(T\cdot C^{2}\cdot H\cdot W)$, which is infeasible for real-time online compensation. In contrast, the simplified formulation preserves the accuracy of accumulated-error estimation (with deviation ¡ 1%), while remaining practical for deployment in real systems.
@@ Appendix C Empirical Study of Error Accumulation Steps
In this section, we extend TCEC to other solvers to demonstrate its generality, such as the most commonly used high-order solver DPM++ (lu2022dpm). The iterative update for DPM-Solver++ (2nd-order variant) is given by:
@@ Appendix F Ablation experiment of regularization term $\lambda_{1}$
Appendix G The impact of inference steps on TCEC performance
Table 1 presents the performance improvements on SDXL-Turbo (4-step), showing that TCEC is likewise applicable to high-speed generative models. However, it is important to note that the gains on SDXL-Turbo are smaller than those on SDXL (50-step). The core reason is that the primary value of TCEC lies in mitigating the accumulation of quantization errors during iterative inference, and thus its performance gains scale directly with the amount of accumulated error. The relatively modest improvements on SDXL-Turbo (4-step) stem from the fact that the small number of steps prevents errors from forming substantial accumulation. Therefore, taking SDXL-Turbo as the baseline, we further conducted additional experiments to investigate the minimal effective number of steps for TCEC.
Table 7: Comparison of SVDQuant and SVDQuant+TCEC across models and steps on MJHQ dataset.
[…]
SDXL 50 SVDQuant 16.6 26.4 - - -
SVDQuant+TCEC 16.0 27.3 0.6 0.9 1.5
SDXL-Turbo 4 SVDQuant 24.3 24.0 - - -
SVDQuant+TCEC 24.5 24.9 -0.2 0.9 0.7
SDXL-Turbo 3 SVDQuant 25.8 18.6 - - -
SVDQuant+TCEC 25.6 18.9 0.2 0.3 0.5
SDXL-Turbo 2 SVDQuant 27.7 17.1 - - -
SVDQuant+TCEC 27.7 17.3 0 0.2 0.2
As shown in Table 7, when the number of iterative steps is $\geq$ 3, TCEC produces meaningful improvements (with gains $\geq$ 0.5 on key metrics). When the number of steps is << 3, the improvements are constrained by the limited amount of accumulated error and are typically << 0.2 (within the range of experimental noise).
Appendix H Inference Overhead of TCEC
@@
TCEC improves performance by performing error correction on the output of the quantized model, and it is completely orthogonal to the backbone quantization algorithm. As shown in Eq. 14, the computation of single-step quantization error involves no complex operations and is accomplished solely through a non-linear mapping: $\varepsilon_{t}=\mathbf{K}_{t}\odot\tilde{\mu}_{\theta}(\tilde{x}_{t},t)$, where $\mathbf{K}_{t}$ denotes a timestep-conditioned channel-wise scaling matrix, and $\odot$ represents element-wise multiplication. Consequently, the additional theoretical computational complexity introduced at each step is $NC^{2}HW$, which is negligible compared to that of the DIT/Unet-Backbone. Practical test data on the Nvidia A800-40G platform show in Table 3 that the extra end-to-end latency incurred is less than 0.5%, with specific test data provided in the table below.
Appendix I Calibration Dataset Ablation
To further verify the impact of the calibration dataset on TCEC performance, we provide an explanation from both theoretical and experimental perspectives.
  • Rationale for selecting the calibration dataset: The size of the calibration set is chosen based on a trade-off between performance and efficiency. As show in Table 8, using SVDQuant-W8A8 as the base quantization algorithm, with COCO as the calibration set and MJHQ as the evaluation set, the results shown in the table indicate that when the calibration set contains fewer than 1024 samples, increasing the number of samples leads to significant performance improvements. However, when the calibration set exceeds 1024 samples and reaches 2048, performance gains plateau. Excessive calibration samples provide no additional benefit and instead prolong the calibration process. Therefore, we select 1024 samples as the standard calibration set size in image generation task.
  • TCEC is insensitive to the domain of the calibration dataset: The input to TCEC’s error-calibration module is the output of the DIT model, which has already undergone the following processing pipeline: 3D-VAE preprocessing → 3D-VAE Encoder inference → DiT inference. The 3D-VAE preprocessing standardizes the input-output distribution, and the subsequent VAE Encoder and DIT inference further reinforce this effect. As a result, the output distributions of the DIT model converge across different datasets, making the error-correction coefficients largely insensitive to the choice of calibration data. In Table 2, OpenSora uses WebVid as the calibration set and VBench—a dataset with markedly different scenes—for evaluation. Despite the domain shift, TCEC consistently improves performance (e.g., +2.08 on W8A8 Imaging Quality and +3.90 on W4A4), providing direct evidence of its robustness to distributional differences.
Table 8: Performance comparison of SDXL with SVDQuant and TCEC at different data sizes.
[…]
SVDQuant - 16.6 26.4 - - -
SVDQuant+TCEC 2048 16.0 27.4 0.6 1.0 1.6
SVDQuant+TCEC 1024 16.0 27.3 0.6 0.9 1.5
SVDQuant+TCEC 512 16.2 26.8 0.4 0.4 0.8
SVDQuant+TCEC 256 16.4 26.7 0.2 0.3 0.5
Appendix J Additional comparison with QNCD
Since TAC (timestep) has not yet released its code, while QNCD (chu2024qncd) has made its code and full experimental details publicly available. We have added comparisons between TCEC and QNCD in our revision, strictly aligning our setup with Table 1 of the QNCD paper.
For 4-bit weight quantization, we adopt BRECQ and AdaRound to preserve model performance. For each experiment, we report the widely used Fréchet Inception Distance (FID) and sFID, and we additionally provide Inception Score (IS) to ensure consistency with previously reported results. All quantization outcomes are obtained on CIFAR-10 (32×32) using a DDIM sampling configuration with 50,000 generated sample:
Table 9: Comparison of TCEC with the error compensation method QNCD.
Method Bitwidth (W/A) DDIM(steps=100) DDIM(steps=250)
[…]
Q-Diffusion 4/8 9.41 4.92 5.13 9.64 4.37 4.59
QNCD 4/8 9.53 4.85 5.06 9.78 4.43 4.51
TCEC 4/8 9.65 4.79 4.91 9.92 4.40 4.45
Q-Diffusion 4/6 7.53 39.07 43.36 7.81 34.65 37.29
QNCD 4/6 8.86 12.26 14.83 9.01 11.09 13.46
TCEC 4/6 9.25 10.76 12.25 10.11 9.67 11.75
The experimental results show that TCEC, which performs cumulative error correction based on theoretically derived error-propagation equations, has a clear advantage over methods that only perform single-step error correction. Under the W4A8, DDIM-100 setting, TCEC improves IS by 0.12 and reduces FID by 0.06 compared with QNCD. Under the lower-activation-bitwidth W4A6 configuration, this advantage becomes even more pronounced: IS improves by 0.39 and FID decreases by 1.50. This is primarily because lower-bit activation quantization leads to significantly amplified cumulative errors, whereas TCEC’s cumulative-error-correction mechanism effectively mitigates this issue.
Appendix K Visual Quality Results.